OpenStudy (anonymous):

Solve the system of equations. 4x - 4y = 10 3x + 2y = 5

OpenStudy (anonymous):

4x - 4y = 10 3x + 2y = 5 we multiply the 2nd one by two 2(3x + 2y = 5) 6x + 4y = 10 4x - 4y - 6x - 4y = 0 -2x -8y = 0 -2x = 8y x = -4y 4(-4y) -4y = 10 -20y = 10 y = - 10/20 y = -1/2 4x - 4(-1/2) = 10 4x - 2 = 10 4x = 12 x = 3 so : x = 3 and y = -1/2 hope that was helpful